4.7g of ahydrated sample of copper 2sphate cuso4.xh20 was heated to a constant mass the mass reduced by 1.71 g what is the value of x
1
Expert's answer
2020-06-15T14:42:00-0400

CuSO4*xH2O (s) + heat --> CuSO4 (s)+ xH2O (g)


The mass of water that has separated is given and equals 1.71 g.

Therefore, the mass of remaining pure CuSO4 is 4.7 - 1.71 = 2.99 g.

n(H2O) = 1.71g / 18g/mol = 0.095 mol

n(CuSO4) = 2.99g / 160g/mol = 0.019 mol

The H2O : CuSO4 mole ratio is 0.095 : 0.019 = 5 : 1.

So the formula is CuSO4*5H2O, and x = 5.

Answer: x = 5